OpenStudy (anonymous):

integral of 1/square root of 16+4x-2x^2 times dx This is a calc 2 question think its trig substitution helppppp

OpenStudy (anonymous):

\[\int\limits_{}^{}1/\sqrt{16+4x-2x^2}\]

OpenStudy (anonymous):

i made the bottom as \[\sqrt{-2(x-1)^2+1}\] then u=x-1 so gotta be \[\int\limits_{}^{}1/\sqrt{18-2u^2}\]

OpenStudy (anonymous):

oh god thanx prony

OpenStudy (anonymous):

*rony

OpenStudy (anonymous):

np =)